x-23 is equal to isotopic mass of 23.02 and x-25 is equal to isotopic mass of 25.147. What is the atomic weight of element x?
1
Expert's answer
2020-09-21T06:35:53-0400

X-23 has isotopic mass of 23.02

X-25 has isotopic mass of 25.147

23/(23+25)=0.479

25/48=0.52

0.479 is the percentage abundance of isotope X-23 while 0.52 is the percentage abundance of isotope X-25

The atomic weight therefore is:

X= (23 x 0.479)+(25 x 0.52)

X= 11.017+13

X=24.017